OpenStudy (anonymous):

Two trains left the station at the same time, one heading due east at 40 miles per hour and the other heading due west at 50 miles per hour. How far apart were they after 4 hours? A. 160 miles B. 200 miles C. 360 miles D. 400 miles WILL GIVE MEDAL

OpenStudy (anonymous):

Relative speed v =40+50 =90 m/hr distance between them=90*4=?

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

can you say that a little more understandable

OpenStudy (anonymous):

90*4=360 miles

OpenStudy (anonymous):

O.K. Iget it

OpenStudy (anonymous):



OpenStudy (anonymous):

here you go

OpenStudy (anonymous):

in one hour they are 90 miles apart.

OpenStudy (anonymous):

O.K.

OpenStudy (anonymous):

can you help me with some more @surjithayer

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

O.K.

OpenStudy (anonymous):

Which information is not necessary to solve this problem? Juanita reads 20 pages per night in her library book. She reads 4 nights a week and her book has 200 pages. How many pages does she read each week? A. Her book has 200 pages. B. She reads 20 pages per night. C.She reads 4 nights a week

OpenStudy (anonymous):

@YoungStudier

OpenStudy (youngstudier):

What do you think?

OpenStudy (anonymous):

I think it is b

OpenStudy (youngstudier):

Um, no, what else?

OpenStudy (anonymous):

c?

OpenStudy (youngstudier):

Ummm, no, not quite, you need both of those things to figure out the question.

OpenStudy (anonymous):

sorry I am bad at math but lititure I am good at

OpenStudy (youngstudier):

So, since it's the only answer choice left, the correct answer is A.

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

did you get my mesege

OpenStudy (youngstudier):

Did you read the pm I sent you? It might help you better grasp what the answer is.